OpenStudy (alex_mattucci):

A company distributes free crayon boxes to all the students of x schools. Each school has (x + 5) classes. The number of students in each class is 2 less than the number of classes in each school. Each student is given 2 boxes of crayons. Part A: Write an expression to show the total number of crayon boxes distributed by the company in x schools. (4 points) Part B: What does x(x + 5) represent? When simplified, what would be the degree and classification of this expression? (4 points) Part C: How can you calculate the total number of students in each school? (2 points)

OpenStudy (alex_mattucci):

i have no clue where to even start

OpenStudy (anonymous):

Part A: Number of schools = x Number of classes in each school = (x + 5) Number of students in each class = Number of classes in each school - 2 = x + 5 - 2 = (x + 3) Number of crayons for each student = 2 Total number of crayons distributed = Number of crayons for each student * Number of students in each class * Number of classes in each school * Number of schools

OpenStudy (anonymous):

hope this helps

OpenStudy (alex_mattucci):

Yes thank you, this helps alot!

OpenStudy (anonymous):

:)